OpenStudy (anonymous):

Math please help! What is the value of the expression 3 – a – b when a = 15 and b = –3? A. –21 B. –15 C. –9 D. 21

OpenStudy (anonymous):

3-15+3=6-15=-9

OpenStudy (anonymous):

thank you!